### v2

# Differentially Private Covariate Balancing Causal Inference

Abstract
Differential privacy is the leading mathematical framework for privacy protection, providing a probabilistic guarantee that safeguards individuals’ private information when publishing statistics from a dataset. This guarantee is achieved by applying a randomized algorithm to the original data, which introduces unique challenges in data analysis by distorting inherent patterns. In particular, causal inference using observational data while preserving privacy is challenging because it requires a good covariate balance between treatment groups, but checking the covariate balance is challenging when privacy is a primary concern because releasing any statistic (e.g., t-statistic) for the balance check compromises privacy. Additionally, the performance of the privatized estimator critically depends on the choice of differential privacy mechanisms, and it remains unexplored how privacy-protecting causal inference could be conducted with observational data. In this article, we present a differentially private two-stage covariate balancing weighting estimator to infer causal effects from observational data. Our algorithm produces both point and interval estimators with statistical guarantees, such as consistency and rate optimality, under a given privacy budget.
Keywords: Covariate balancing propensity score, Weighted average treatment effect, Empirical risk minimization, Differential privacy.
1 Introduction
In the digital era, the secure handling of sensitive data has become essential for research, policy-making, and business. As data collection expands across various sectors, the risk of compromising individual privacy increases. Despite these concerns, sensitive data remains critical for informed, evidence-based decision-making, necessitating methods that balance data accessibility with privacy protection. Differential privacy (DP) has emerged as the gold standard for achieving this balance. It is increasingly adopted in industry (Erlingsson et al., 2014; Apple, 2017) and by government agencies, such as the U.S. Census Bureau (Abowd, 2018). DP offers a probabilistic guarantee of privacy, safeguarding against arbitrary breaches by applying a privacy mechanism (e.g., adding random noise) to summary statistics and synthetic data before their release to the public. However, while DP effectively protects privacy, integrating it into statistical analyses introduces significant challenges.
In particular, causal inference is essential for decision-making across various fields. Randomized experiments are ideal for identifying causal effects, but their costs and complexities often necessitate the use of observational data. However, observational studies face challenges, particularly in adjusting for confounding variables, which can bias treatment effect estimates. The propensity score (Rosenbaum and Rubin, 1983), the probability of treatment given covariates, helps eliminate confounding bias when treatment relies on observables and has been adopted in many applications like matching, stratification, and weighting (Rosenbaum and Rubin, 1985; Abadie and Imbens, 2006; Heckman et al., 1997). Rosenbaum and Rubin (1983) demonstrated that the true propensity score mitigates the confounding bias by creating quasi-randomized experiments in which covariate distributions are well-balanced between treatment groups. However, accurately specifying the true propensity score model is typically difficult, and even slight misspecifications can introduce significant bias (Kang and Schafer, 2007).
To ensure robust analyses under the misspecification of the propensity score, the focus of the propensity score analysis literature has shifted from accurately predicting treatment assignment to instead using it to achieve a good covariate balance between treatment groups (Stuart, 2010; Imai and Ratkovic, 2014; Imbens and Rubin, 2015). In this regard, the covariate balance check is an essential step in enhancing the credibility of any propensity score analysis. However, checking the covariate balance is challenging when privacy is a primary concern because releasing any statistic (e.g., t-statistic) for the balance check compromises privacy, and their secure handling incurs an additional privacy cost. Additionally, as the balance check is typically an iterative process, each iteration increases the risk of information leakage, posing unique challenges for performing robust inference by maintaining the covariate balance while ensuring privacy protection. A promising approach is the covariate balancing weighting scheme (Hainmueller, 2012; Chan et al., 2015; Zhao and Percival, 2017; Zhao, 2019; Hazlett, 2020; Kong et al., 2023; Fan et al., 2023; Huling and Mak, 2024). These methods ensure automatic covariate balance by solving optimization problems for propensity score estimation under empirical covariate balancing constraints. However, these methods have only been developed in non-private settings; research on covariate balancing inference in a private setting remains unexplored.
In light of the unique challenges associated with robust causal analysis in privacy-sensitive contexts, this article introduces a privacy-preserving, covariate-balancing causal inference methodology. Our work contributes to the current body of literature in several ways. First, we discuss privatization strategies and the selection of privacy mechanisms to achieve desirable asymptotic properties, such as consistency and rate optimality. We then propose a two-stage privatization algorithm within a unified framework for estimating a wide range of causal effects from observational data. Leveraging the covariate-balancing scoring rules (CBSR) of Zhao (2019), we obtain weights that are robust to propensity-score misspecification in privacy-sensitive settings. This choice of balancing framework is pivotal for selecting a compatible privacy mechanism, as its convex, differentiable objective pairs naturally with DP methods that privatize gradients. We instantiate privacy with the K-Norm Gradient (KNG) mechanism (Reimherr and Awan, 2019), which releases a noise-perturbed score and solves the resulting estimating equations, thereby preserving CBSR’s balance properties while ensuring privacy. By contrast, generic mechanisms such as the exponential mechanism (McSherry and Talwar, 2007) and objective/output perturbation (Chaudhuri et al., 2011) do not, in general, maintain the required balance constraints or utility guarantees for weighting, and thus may fail to deliver the desired properties. Additionally, with appropriate choices of the covariate balancing framework and privacy mechanism, our privatized estimator exhibits favorable asymptotic properties, such as consistency, rate optimality, and asymptotic covariate balance, while preserving privacy. We also provide asymptotically valid confidence intervals for the estimated causal effects. Through comprehensive simulation studies across various privacy budgets and sample sizes, we evaluate the performance of our methodology and compare it against existing DP causal inference approaches. The results indicate that our methodology exhibits robust performance under both correctly specified and misspecified propensity score models. Finally, we apply our methodology to real-world data from a job training program evaluation, successfully recovering the non-private estimates and achieving satisfactory covariate balance. This application underscores the practical utility and effectiveness of our approach in privacy-sensitive contexts. All technical proofs of theorems and lemmas are provided in the supplementary material.
The rest of the paper is organized as follows. Section 2 presents the preliminaries for the differential privacy and covariate balancing causal inference framework. Section 3 provides our alrorithm and proved its privacy guarantees and statistical properties. Section 4 provides simulation studies for validating our methodology developed in the previous sections, and Section 5 provides an application of our methodologies to real-world data of a job training program. Section 6 concludes with some final discussion.
1.1 Current literature
Although DP is a rapidly expanding field, research focusing on propensity score analyses for observational data in privacy-sensitive contexts, similar to ours, remains limited. Lee et al. (2019) proposed a privacy-preserving inverse propensity score estimator for estimating the average treatment effect (ATE). They suggested a Horvitz-Thompson-type estimator using an objective perturbation technique to ensure privacy. However, this approach requires regularization, resulting in a residual bias in the propensity score, even asymptotically. Additionally, they did not consider private confidence intervals or explore covariate balance, which is central to causal inference from observational data. Guha and Reiter (2024) developed a causal inference methodology that leverages the subsample-and-aggregate algorithm (Nissim et al., 2007) to estimate the weighted average treatment effects with binary outcomes. They also presented private standard errors and confidence intervals for the estimator. However, they did not consider the covariate balance that we consider in this paper, and the performance guarantees of their estimator depend on the number of subgroups, which is a hyperparameter for the algorithm that analysts must tune for the application at hand.
Some authors have addressed causal inference problems under different DP paradigms or study designs for causal inference than those used in our work. D’Orazio et al. (2015) introduced differential privacy mechanisms in causal inference and algorithms for releasing private estimates of causal effects, mainly for experimental settings. Kusner et al. (2016) explored causal inference using the additive noise model, a more restrictive approach than the potential outcomes framework considered in this paper. Komarova and Nekipelov (2020) demonstrated that under differential privacy, identifying causal parameters fails in regression discontinuity designs. Agarwal and Singh (2021) and Ohnishi and Awan (2025) proposed causal inference methods under the local DP model, and Niu et al. (2022) introduced a meta-algorithm for privately estimating conditional average treatment effects, without addressing private confidence intervals or covariate balance. Javanmard et al. (2024) proposed Cluster-DP for randomized trials, while Chen et al. (2024) studied the experimental design problem under Distributed DP with secure aggregation.
2 Preliminaries
2.1 Notation and Causal Estimands
Throughout this manuscript, we adopt the Rubin Causal Model (Imbens and Rubin, 2015) as our causal inference framework. We consider $n$ units, indexed by $i=1,\ldots,n$, as a random sample from a large super-population. Each unit $i$ has an outcome $Y_{i}\in[0,1]$, treatment assignment $Z_{i}\in\{0,1\}$, and covariates $X_{i}\in\mathcal{X}$, respectively. We will assume throughout that $\mathcal{X}$ is the unit ball so that $\|X_{i}\|_{2}\leq 1$. This boundedness assumption is standard practice in the differential privacy literature (e.g., Lei et al. (2017), Ferrando et al. (2022),Chaudhuri et al. (2011) to name a few). We consider a binary treatment with the unknown assignment mechanism $e(x)=p(Z_{i}=1\mid X_{i}=x)$, which we call propensity score. We assume that there is neither interference nor hidden versions of treatment, $Y_{i}(z)$ denote a potential outcome for $Z_{i}=z$. We make a common set of assumptions that enable us to identify causal effects.
Assumption 1 (Positivity).
There exists a positive constant $0<\eta\leq 0.5$, such that the probability of treatment assignment given the covariates is bounded as $\eta\leq e(X)\leq 1-\eta$ for $X\in\mathcal{X}$.
Assumption 2 (Unconfoundedness).
The potential outcomes are conditionally independent of treatment assignment given the covariates: $\{Y_{i}(0),Y_{i}(1)\}\perp\!\!\!\perp Z_{i}\mid X_{i}$.
We define the average treatment effect (ATE) as $\tau=\mathrm{E}\{Y(1)-Y(0)\}$. Assuming the existence of the marginal density function for the covariates $X$, denoted as $f(x)$, with respect to a base measure $\mu$, we can write the ATE as $\tau=\int\tau(x)f(x)dx$, where $\tau(x)=\mathrm{E}\{Y(1)-Y(0)\mid X=x\}$. The ATE is widely valued due to its ability to provide interpretable estimates of treatment effects on the whole population. However, in practice, the exclusive focus on the ATE is a significant limitation for several reasons. First, the ATE may reflect the effect of an intervention that is practically infeasible or impossible to apply to every unit in the study population. Second, the available data may not accurately capture the characteristics of the intended population of interest. In such cases, conventional analysis methods may fail to yield an estimate of the average treatment effect for the appropriate target population. Finally, researchers frequently exclude extreme or atypical units from their analysis, resulting in estimates that pertain only to a subpopulation within the original target population. A recent body of literature suggests focusing on subpopulations exhibiting sufficient covariate overlap between treatment groups. Li et al. (2018) introduced a class of balancing weights designed to balance the distributions of covariates between comparison groups for any predetermined target population. They consider the target population density, expressed as the product of the marginal density $f(x)$ and a predefined function $h(x)$ of the covariate $x$. Within this framework, a general class of estimands is defined as the weighted average treatment effect (WATE) over the target population:
$\tau_{h}=\frac{\int\tau(dx)f(x)h(x)\mu(dx)}{\int f(x)h(x)\mu(dx)},$ (1)
where $h(x)$ is a known function of the covariates. By choosing different forms of the function $h(\cdot)$, the WATE can capture various estimands of causal effects: $h(x)=1$ for ATE, $h(x)=e(x)$ for the ATE on the treated (ATT), h(x)=1e(x)h(x)=1-e(x) for the ATE on the control (ATC).
2.2 Covariate Balancing Propensity Score Estimation
Generally, propensity score analysis can be viewed as a decision problem through the lens of statistical decision theory. Our primary goal for propensity score estimation is to select an element $P$ as the prediction from $\mathcal{P}$, a convex class of (conditional) probability measures on some general sample space $\Omega$. The prediction is evaluated by the scoring rule, an extended real-valued function $S:\mathcal{P}\times\Omega\rightarrow[-\infty,\infty]$ such that $S(P,\cdot)$ is integrable for all $P\in\mathcal{P}$ (Gneiting and Raftery, 2007). If the decision is $P$ and $\omega$ is a realization, the utility (or loss function) is written as $S(P,\omega)$. If the outcome is probabilistic and the actual probability distribution is $Q$, the expected score of predicting $P$ is $S(P,Q)=\int S(P,\omega)Q(dw)$. A scoring rule $S$ is said to be proper if $S(Q,Q)\geq S(P,Q),\forall P,Q\in\mathcal{P}$, and strictly proper if the equality holds only when $P=Q$.
Given a strictly proper scoring rule $S$, the maximum score estimator of $\theta$ is obtained by maximizing the average score. Given observations $D=\{X_{i},Y_{i},Z_{i}\}_{i=1}^{n}$,
$\hat{\theta}_{n}=\arg\max_{\theta}S_{n}(\theta,D)=\arg\max_{\theta}\frac{1}{n}\sum_{i=1}^{n}S\{e_{\theta}(X_{i}),Z_{i}\}.$ (2)
If $S$ is differentiable, the maximizer of $\mathrm{E}\{S_{n}(\theta)\}$ (the population version of (2)) satisfies the estimating equations: $\nabla_{\theta}\mathrm{E}\{S_{n}(\theta,D)\}=0$.
In observational studies with binary treatment ($\Omega=\{0,1\}$), a probability distribution $P$ can be characterized by an assignment probability $0\leq e\leq 1$. Savage (1971) showed that every real-valued proper scoring rule $S$ can be written as $S(e,1)=G(e)+(1-e)G^{\prime}(e)$, $S(e,0)=G(e)-eG^{\prime}(e)$, where $G:[0,1]\to\mathbb{R}$ is a convex function. If $G$ is second-order differentiable, this can be represented as:
$\frac{\partial}{\partial e}S(e,z)=(z-e)G^{\prime\prime}(e)\text{ for }z=0,1.$ (3)
While many choices exist for $G$, we consider the Beta family for the class of proper scoring rules:
$G^{\prime\prime}_{\alpha,\beta}(e)=e^{\alpha-1}(1-e)^{\beta-1},\quad-\infty<\alpha,\beta<\infty.$ (4)
Zhao (2019) proposed the covariate balancing scoring rule (CBSR) using the Beta family (4), interpreting the estimating equation as the first-order covariate balancing constraint. Suppose we have i.i.d. observations $D_{i}=(X_{i},Y_{i},Z_{i})$ for $i=1,\ldots,n$, and we fit a model for the propensity score in a family $\mathcal{P}=\{e_{\theta}(X):\theta\in\Theta\}$. We estimate $e_{\theta}(X)$ with the sieve logistic regression model (Geman and Hwang, 1982), where we adopt an orthogonalized polynomial series of the covariates for the propensity score estimation. The sieve estimator is a class of non-parametric estimators that use progressively more complex models, such as higher-order moment predictors, to estimate an unknown high-dimensional function as more data becomes available. The details of the sieve estimator are provided in the supplementary material. Consider the logistic link function with finite-dimensional regressors $\phi(X)=(\phi_{1}(X),\ldots,\phi_{d}(X))^{\top}$, where $\phi_{j}:\mathcal{X}\to\Phi$ with $\|\phi(X)\|_{2}\leq C_{\phi}$ for some constant $C_{\phi}\in(0,\infty)$: $e_{\theta}(X)=l^{-1}\{g_{\theta}(X)\}=l^{-1}\{\theta^{\top}\phi(X)\}$, where $l$ is the logistic link function: $l(e)=\log\left(\frac{e}{1-e}\right)$, $l^{-1}(g)=\frac{\exp(g)}{1+\exp(g)}$. Using representation (3) and the inverse function theorem, we can rewrite the estimating equation as:
$\nabla_{\theta}\mathrm{E}\{S_{n}(\theta,D)\}=\mathrm{E}(\nabla_{\theta}S[l^{-1}\{\theta^{\top}\phi(X)\},Z])=\mathrm{E}[\{Z-(1-Z)\}w_{\theta}(X,Z)\phi(X)]=0,$ (5)
where $w_{\theta}(x,z)=\frac{G^{\prime\prime}\{e_{\theta}(x)\}}{l^{\prime}\{e_{\theta}(x)\}}[z\{1-e_{\theta}(x)\}+(1-z)e_{\theta}(x)]$. This is exactly the first-order balancing constraint of Imai and Ratkovic (2014) with the weighting function $w_{\theta}(x,z)$ determined by the scoring rule and link function. The maximum score estimator $\hat{\theta}_{n}$ can be obtained by solving these equations empirically, implying that $\hat{\theta}_{n}$ automatically achieves the empirical first-order covariate balance, $\sum_{i=1}^{n}Z_{i}w(X_{i},Z_{i})\phi(X_{i})=\sum_{i=1}^{n}(1-Z_{i})w(X_{i},Z_{i})\phi(X_{i})$.
Notably, Zhao (2019) also showed that using the Beta family scoring rule (4), the most commonly used WATE can be expressed as the weighted average treatment effects with $h_{\alpha,\beta}=e_{\theta}(X)^{\alpha+1}\{1-e_{\theta}(X)\}^{\beta+1}$ with specific values of $\alpha$, $\beta$ $\in[-1,0]$. We can rewrite (1) as:
$\tau_{\alpha,\beta}=\frac{\mathrm{E}[h_{\alpha,\beta}(X)\{Y(1)-Y(0)\}]}{\mathrm{E}\{h_{\alpha,\beta}(X)\}},$ (6)
and estimate it by
$\hat{\tau}_{\alpha,\beta}=\frac{\sum_{i=1}^{n}Z_{i}w_{\hat{\theta}_{\alpha,\beta}}(X_{i},1)Y_{i}}{\sum_{i=1}^{n}Z_{i}w_{\hat{\theta}_{\alpha,\beta}}(X_{i},1)}-\frac{\sum_{i=1}^{n}(1-Z_{i})w_{\hat{\theta}_{\alpha,\beta}}(X_{i},0)Y_{i}}{\sum_{i=1}^{n}(1-Z_{i})w_{\hat{\theta}_{\alpha,\beta}}(X_{i},0)},$ (7)
with $\hat{\theta}_{\alpha,\beta}$ being estimated with corresponding score functions for the specific values of $\alpha$ and $\beta$. Table 1 summarizes the discussion in this section about the correspondence of estimands, sample weighting functions, and score functions. Note that our methodology starts with choosing the causal quantity of interest and corresponding values for $\alpha$ and $\beta$, through which the score function is automatically defined.
Table 1: Correspondence of estimands, sample weighting functions, and the score functions for different values of $\alpha$ and $\beta$. ATO represents the average treatment effect on the overlapped population Li et al. (2018).
$\alpha$ $\beta$ $\tau_{\alpha,\beta}$ $w(x,1)$ $w(x,0)$ $S(e,1)$ $S(e,0)$ -1 -1 ATE $\frac{1}{e(x)}$ $\frac{1}{1-e(x)}$ $\log\frac{e}{1-e}-\frac{1}{e}$ $\log\frac{1-e}{e}-\frac{1}{1-e}$ -1 0 ATC $\frac{1-e(x)}{e(x)}$ 11 $\log\frac{1-e}{e}$ $\frac{1}{e}$ 0 -1 ATT $\frac{1}{e(x)}$ $\frac{1}{1-e(x)}$ $\log\frac{e}{1-e}$ $\frac{1}{1-e}$ 0 0 ATO 1e(x)1-e(x) $e(x)$ $\log e$ $\log(1-e)$
Remark (Synthesis with privacy mechanism).
Among covariate-balancing propensity score methods, entropy balancing (Hainmueller, 2012) is widely used. It reweights units by solving a maximum-entropy program that enforces prespecified covariate moments to match across groups, yielding a balanced pseudo-population under ignorability. Chan et al. (2015) propose empirical balancing calibration weighting, which constructs calibration weights by minimizing a divergence from uniform subject to linear balance constraints on chosen basis functions. However, both approaches are ill-suited for privacy-preserving weighting: standard privacy mechanisms require bounded losses or gradients to calibrate noise, whereas the entropy and calibration objectives generally violate this boundedness. By contrast, the CSBR framework satisfies the required boundedness and is therefore compatible with private weighting.
2.3 Differential Privacy
We consider the central DP model, which involves a trusted data curator collecting and storing sensitive data from individuals in a central location. The data curator then performs a privacy-preserving analysis of the data and releases the results to the public. Let $\mathcal{D}_{n}$ denote the collection of databases with $n$ units. Let $\mathcal{M}$ be a randomized algorithm that takes a database $D\in\mathcal{D}_{n}$ as input and outputs a random quantity $r$, i.e., $\mathcal{M}(D)=r$. We say $D,D^{\prime}\in\mathcal{D}_{n}$ are adjacent if $d_{\mathrm{Ham}}(D,D^{\prime})=1$, where $d_{\mathrm{Ham}}(D,D^{\prime})$ is the Hamming distance between $D$ and $D^{\prime}$.
Definition 1 ($\epsilon$-Differential Privacy).
An algorithm $\mathcal{M}$ satisfies $\epsilon$-differential privacy ($\epsilon$-DP), if for any pair of adjacent databases $D,D^{\prime}\in\mathcal{D}_{n}$, and any measurable set $S\subseteq\text{range}(\mathcal{M})$, $\mathrm{pr}\{\mathcal{M}(D)\in S\}\leq\exp(\epsilon)\mathrm{pr}\{\mathcal{M}(D^{\prime})\in S\}.$
The definition states that $\mathcal{M}$ satisfies $\epsilon$-DP when the distributions of its outputs are similar for any two adjacent databases, where $\epsilon$ measures the similarity. Intuitively, if a record in the database changed from $d$ to $d^{\prime}$, the output distribution of $M$ would be similar, making it difficult for an adversary to determine whether any record is present in the database or not. The value $\epsilon$ is called the privacy budget, and lower values correspond to a stronger privacy guarantee.
Two important properties of differential privacy are composition and invariance to post-processing (Dwork and Roth, 2014). Composition allows one to derive the cumulative privacy cost when releasing the results of multiple privacy mechanisms: if $\mathcal{M}_{1}$ is $\epsilon_{1}$-DP and $\mathcal{M}_{2}$ is $\epsilon_{2}$-DP, then the joint release $(\mathcal{M}_{1}(D),\mathcal{M}_{2}(D))$ satisfies $(\epsilon_{1}+\epsilon_{2})$-DP. Invariance to post-processing ensures that applying a data-independent procedure to the output of a DP mechanism does not compromise the privacy guarantee: if $\mathcal{M}$ is $\epsilon$-DP with range $\mathcal{Y}$, and $f:\mathcal{Y}\rightarrow\mathcal{Z}$ is a (potentially randomized) function, then $f\circ\mathcal{M}$ is also $\epsilon$-DP.
[…]
Proposition 1 (Laplace Mechanism).
Let $f:\mathcal{D}_{n}\to\mathbb{R}^{k}$. The Laplace mechanism is defined as $M(D)=f(D)+(\nu_{1},...,\nu_{k})^{\top},$ where the $\nu_{i}$ are independent Laplace random variables, $\nu_{i}\sim\mathrm{Lap}(0,\Delta f/\epsilon)$, where the density of the Laplace distribution, $\mathrm{Lap}(\mu,b)$, is $f(\nu|\mu,b)=\frac{1}{2b}\exp(-\frac{|\nu-\mu|}{b})$. Then $M$ satisfies $\epsilon$-DP.
Reimherr and Awan (2019) introduced the K-Norm Gradient Mechanism (KNG). This mechanism is especially useful for problems involving the minimization of an objective function through which the parameters of interest are obtained.
Proposition 2 (K-Norm Gradient Mechanism (KNG)).
Let $\Theta\subset\mathbb{R}^{d}$ be a convex set, $\|\cdot\|_{K}$ be a norm on $\mathbb{R}^{d}$, and $\nu$ be the Lebesgue measure on $\Theta$. Let $\{\ell_{n}(\theta;D):\Theta\rightarrow\mathbb{R}\mid D\in\mathcal{D}_{n}\}$ be a collection of measurable functions whose gradient is defined almost everywhere. We say that this collection has sensitivity $\Delta:\Theta\rightarrow\mathbb{R}_{+}$, if $\|\nabla\ell_{n}(\theta;D)-\nabla\ell_{n}(\theta;D^{\prime})\|_{K}\leq\Delta(\theta)<\infty$, for all adjacent $D,D^{\prime}\in\mathcal{D}_{n}$ and $\theta$. If $\int_{\Theta}\exp\left\{-\frac{1}{\Delta(\theta)}\|\nabla\ell_{n}(\theta;D)\|_{K}\right\}d\nu(\theta)<\infty$ for all $D\in\mathcal{D}_{n}$, then the collection of probability measures $\{\mu_{D}\mid D\in\mathcal{D}_{n}\}$ with densities (with respect to $\nu$) given by
$f_{D}(\theta)\propto\exp\left\{-\frac{\epsilon}{2\Delta(\theta)}\|\nabla\ell_{n}(\theta;D)\|_{K}\right\}$ (8)
satisfies $\epsilon$-DP.
The KNG method starts with an objective function and favors summaries that nearly minimize it by weighting according to how close the gradient is to zero. Under certain technical conditions (Reimherr and Awan, 2019, Theorem 3.2), the KNG achieves an asymptotic error of $O_{p}(n^{-1})$, which is asymptotically negligible compared to the statistical error for many problems. In contrast, under the same conditions, the exponential mechanism introduces noise of magnitude $O_{p}(n^{-1/2})$ (Awan et al., 2019). Furthermore, unlike objective perturbation, KNG does not require regularization.
Remark.
One condition for the KNG to achieve $O_{p}(n^{-1})$ error is that the loss function is strongly convex. Without the strong convexity (only with convexity), the KNG is still a valid $\epsilon$-DP mechanism, which is an additional advantage over other mechanisms that require strong convexity (e.g., objective perturbation), but may have error greater than $O_{p}(n^{-1})$. The negative score function $S\{e_{\theta}(x),z\}$, which we use as the loss function in this study, is a convex function for $z\in\{0,1\}$ and $1\leq\alpha,\beta\leq 1$, and $S$ is strongly convex for all $1\leq\alpha,\beta\leq 1$ except for $\alpha=-1,\beta=0$ and $\alpha=0,\beta=-1$ (Zhao, 2019, Theorem 3.2).
Remark (Choice of privacy mechanism).
Among the methods to achieve DP, the exponential mechanism (McSherry and Talwar, 2007) is a popular mechanism for its flexibility and adaptability across various statistical analyses. This mechanism is especially useful for problems involving the minimization of an objective function $\ell_{n}(\theta,D)$ for $D\in\mathcal{D}_{n}$, which encompasses a wide range of statistical tasks. The exponential mechanism releases an estimate $\tilde{\theta}$ based on the density: $f(\theta)\propto\exp\{-c_{0}\ell_{n}(\theta,D)\},$ where $c_{0}$ is a constant determined by the sensitivity of $\ell_{n}$ and the desired level of privacy. However, as Awan et al. (2019) demonstrated, the noise added by the exponential mechanism can be significant, sometimes exceeding that of other mechanisms, thereby reducing its utility. Another popular mechanism that involves minimizing an objective function is the output/objective perturbation (Chaudhuri et al., 2011), which Lee et al. (2019) adopted for the propensity score estimation. A limitation of these methods for our purposes is their requirement of regularization, which introduces bias that can remain even asymptotically and breaks the covariate balance and leads to inconsistent estimates. Additionally, the regularization parameter is typically determined via cross-validation. Each iteration of cross-validation increases the risk of information leak, thus, it is desirable to define the objective function for the propensity score without regularization to maintain the balance.
3 Methodologies
3.1 Differentially Private Covariate Balancing Estimator
This section introduces a privacy-preserving covariate balancing algorithm for inferring the WATE and provides theoretical guarantees for the estimator. Our goal is to infer the WATE (6) without compromising the privacy guarantee. However, the CBSR estimator (7) involves sensitive information about individuals; therefore, releasing this estimator to the public can lead to serious privacy leakage. A naïve approach to privatizing the estimator is to calculate the estimator’s sensitivity directly and add calibrated noise to the estimator before releasing it. The difficulty of this approach lies in the fact that the estimator involves the weighting function $w(X_{i},z)$. The weighting function is estimated by solving an optimization problem using all $n$ data points in a database $D$. Therefore, the estimated value of $w$ can vary for all units $i=1,\ldots,n$, even when estimated with adjacent databases $D,D^{\prime}\in\mathcal{D}_{n}$ that differ in only one record. When considering the direct privatization of the Hájek-type weighted estimator (7), its naïve sensitivity is $|\tau(D)|\leq 1$, leading to calibrated noise of magnitude $O_{p}(1)$, which dominates the statistical error $O_{p}(n^{-1/2})$. Guha and Reiter (2024) considered the subsample and aggregate algorithm (Nissim et al., 2007) to achieve $\epsilon$-DP, using a suboptimal sensitivity for the estimators within $M$ subsamples. Their algorithm splits $D$ into $M$ disjoint subsets, performs propensity score analysis within each subset, and then aggregates the estimates. They demonstrated that their estimator achieves consistency as the number of observations within each subset approaches infinity. However, the efficiency depends on the number of subsets $M$, and the analysts must choose it in advance to ensure favorable asymptotic properties.
In light of these challenges, our algorithm consists of two privatization steps. The first step is to privatize $\hat{\theta}_{n}$ obtained by approximately solving (2) with the KNG mechanism. The second step is to compute an estimator using $e_{\tilde{\theta}}(X_{i})$ for (7) and release the final privatized estimator $\tilde{\tau}^{(2)}$ by applying the Laplace mechanism to each of four components which make up $\tilde{\tau}^{(2)}$.
3.2 Minimax Risk Lower Bound for WATE Estimation
First, we discuss a lower bound for the WATE estimation problem under the central DP model, where a data curator has access to individual data and then applies a differential privacy mechanism and releases the privatized outputs (e.g., summary statistics) to an untrusted data analyst. According to Barber and Duchi (2014), the minimax lower bound of the mean squared error for 1-dimensional mean estimation under the central model is $O[\{n^{-1}+(\epsilon n)^{-2}\}]$. We let $\mathcal{M}_{\epsilon}$ denote the set of all privacy mechanisms that satisfy $\epsilon$-DP. Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{n}$ are drawn according to some distribution $P\in\mathcal{P}$, where $\mathcal{P}$ denotes a class of distributions on the sample space of covariates, potential outcomes and treatment assignment variables. Also, we define an estimator $\hat{\tau}$ for WATE $\tau$ as a measurable function that maps inputs to a real value, that is, $\hat{\tau}:\Omega^{n}\to\mathbb{R}$, where $\Omega$ generally denotes the space of inputs.
Lemma 1.
There exists a constant $c$ such that
$c\{n^{-1}+(\epsilon n)^{-2}\}\leq\inf_{M_{\epsilon}\in\mathcal{M}_{\epsilon}}\inf_{\hat{\tau}}\sup_{P\in\mathcal{P}}\mathrm{E}\{(\hat{\tau}-\tau)^{2}\}$ (9)
Lemma 1 is a simple modification of Barber and Duchi (2014, Proposition 2), implying that if a DP WATE estimation procedure achieves the minimax lower bound $O\{n^{-1}+(\epsilon n)^{-2}\}$, then it is minimax optimal among all $\epsilon$-DP procedures. Failing to match the bound does not necessarily imply the method is suboptimal as it may be the case that the lower bound is simply not tight.
3.3 Private Propensity Score Estimation
Table 2: Correspondence of estimands and sensitivity for different values of $\alpha$ and $\beta$.
$\alpha$ $\beta$ $\tau_{\alpha,\beta}$ $\Delta_{\theta,\alpha,\beta}$ $\Delta_{V,\alpha,\beta}$ -1 -1 ATE $2C_{\phi}/\eta$ $(2n\eta)^{-1}$ -1 0 ATC $2C_{\phi}(1-\eta)/\eta$ $(2n\eta^{2})^{-1}$ 0 -1 ATT $2C_{\phi}(1-\eta)/\eta$ $(2n\eta^{2})^{-1}$ 0 0 ATO $2C_{\phi}(1-\eta)$ $(2n\eta^{2}(1-\eta))^{-1}$
We first consider estimating the propensity score in a private manner via the KNG mechanism with $\lVert\cdot\rVert_{K}=\lvert\cdot\rVert_{2}$. We consider privatizing the parameter $\hat{\theta}_{n}$ obtained by solving (2). Given a privacy budget $\epsilon$, we draw $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from the density:
$f(\theta)\propto\exp\left\{-\frac{p\epsilon}{2\Delta_{\theta,\alpha,\beta}}\|\nabla S_{n,\alpha,\beta}(\theta;D)\|_{2}\right\},$ (10)
where $p\in(0,1)$, $S_{n,\alpha,\beta}(\theta;D)$ and $\Delta_{\theta,\alpha,\beta}$ denote the fraction of the privacy budget for the propensity score privatization, the loss function, and the $\ell_{2}$-sensitivity of $S_{n,\alpha,\beta}$, respectively. Releasing $\tilde{\theta}^{(1)}_{\alpha,\beta}$ satisfies $p\epsilon$-DP. To implement (10), we need to compute the gradient and sensitivity of the loss $S_{n,\alpha,\beta}(\theta;D)$, which are given by the following lemma.
Lemma 2.
Suppose we use the Beta family in (4). For $\alpha,\beta\in[-1,0]$, we have $\nabla_{\theta}S_{n,\alpha,\beta}(\theta;D)=\sum_{i=1}^{n}\left[\{Z_{i}-e_{\theta}(X_{i})\}e_{\theta}(X_{i})^{\alpha}\{1-e_{\theta}(X_{i})\}^{\beta}\right]\{\phi_{1}(X_{i}),\ldots,\phi_{d}(X_{i})\}^{\top}$ and $\Delta_{\theta,\alpha,\beta}\leq 2C_{\phi}|\{Z_{i}-e_{\theta}(X_{i})\}e_{\theta}(X_{i})^{\alpha}\{1-e_{\theta}(X_{i})\}^{\beta}|.$
By Assumption 1, $\Delta_{\theta,\alpha,\beta}$ is bounded for $\alpha,\beta\in[-1,0]$. Table 2 presents the correspondence of estimands and sensitivity for different combinations of $\alpha$ and $\beta$.
Given the appropriate gradient and sensitivity for each estimand, we sample from (10) to obtain a private estimator $\tilde{\theta}^{(1)}_{\alpha,\beta}$ using the privacy-aware rejection sampler (Awan and Rao, 2024). This sampling technique allows for exact sampling from the target density (10). The details are provided in the supplementary material. We then plug in $\tilde{\theta}^{(1)}_{\alpha,\beta}$ for $\hat{\tau}_{\alpha,\beta}$ in (7) and obtain
$\tilde{\tau}^{(1)}_{\alpha,\beta}=\frac{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)Y_{i}}{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)}-\frac{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)Y_{i}}{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)}.$ (11)
Since we adopt Assumption 1, we truncate $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ to $\eta,1-\eta]$, which is needed to learn the sensitivity in the next section. $\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}$ is obtained by plugging in $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ for the weighting function in Table 1. Note that the estimator (11) does not satisfy $\epsilon$-DP and is vulnerable to adversarial attacks because it contains private information. Thus, we need an additional privatization step, which we develop in the following section. Prior to that, we first present the following theorem.
Theorem 3.1 (Asymptotic covariate balance).
Suppose we use the CBSR framework with the Beta-family scoring rule defined by equations (3) and (4) and the KNG mechanism to privatize the estimator under Assumption 1 and 2. Then, for all $j=1,\ldots,d$, we have the first-order covariate balancing condition, $\delta_{n}(\tilde{\theta}^{(1)}_{\alpha,\beta})=\sum_{i=1}^{n}\{Z_{i}-(1-Z_{i})\}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},Z_{i})\phi_{j}(X_{i})=O_{p}(1/n)$.
This theorem states that our estimator uses a privatized weighting function, $\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}$, to achieve asymptotic covariate balance. This covariate balance enhances both the estimator’s accuracy and the credibility of the analysis, as suggested by existing literature that emphasizes the critical role of covariate balance in ensuring reliable causal inference, especially under model misspecification (Imai and Ratkovic, 2014; Chan et al., 2015; Zhao and Percival, 2017; Fan et al., 2023).
3.4 Private WATE Estimation
The second stage of privatization applies the Laplace mechanism independently to the numerators and denominators of (11). Specifically, we consider the following private estimator:
$\tilde{\tau}^{(2)}_{\alpha,\beta}=\frac{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)Y_{i}+\nu_{1}}{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)+\nu_{2}}-\frac{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)Y_{i}+\nu_{3}}{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)+\nu_{4}},$ (12)
where $\nu_{j}\sim\mathrm{Lap}\{0,1/(1-p)q_{j}\epsilon\eta\}$ and $0<q_{j}<1$ for $j=1,\ldots,4$, and $\sum_{j=1}^{4}q_{j}=1$ and the $\ell_{1}$-sensitivity of each component is $1/\eta$. The quantities $q_{j}$ represent the customizable privacy budget for each component of (11). Algorithm 1 summarizes the process for obtaining the privatized point estimator $\tilde{\tau}^{(2)}_{\alpha,\beta}$. Proposition 3 establishes the privacy guarantee of Algorithm 1.
Algorithm 1 Differentially Private Covariate Balancing Estimator
0: Database $D=\{(X_{i},Z_{i},Y_{i})\}_{i=1}^{n}$, causal estimands of interest in Table 1, the corresponding score function $S_{n,\alpha,\beta}$ and sensitivity $\Delta_{\theta,\alpha,\beta}$, positivity bound $\eta$, privacy budget $\epsilon$, fraction of privacy budget allocation parameters $p$ and $q_{j}$ for $j=1,\ldots,4$.
0: Differentially private estimator $\tilde{\tau}^{(2)}_{\alpha,\beta}$.
1:First Stage: Privatization of Propensity Score
  1. 1.
    Using the privacy-aware rejection sampler (Awan and Rao, 2024), draw the private propensity score parameter $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from:
    $f(\theta)\propto\exp\left\{-\frac{p\epsilon}{2\Delta_{\theta,\alpha,\beta}}\|\nabla S_{n,\alpha,\beta}(\theta;D)\|_{2}\right\}.$
  2. 2.
    Compute the privatized propensity score $e_{\tilde{\theta}^{(1)}}(X_{i})$ using $\tilde{\theta}^{(1)}$.
  3. 3.
    Truncate $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ to $\eta,1-\eta]$.
2:Second Stage: Release of the Final Privatized Estimator
  1. 1.
    Using $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$, calculate the weights $\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},z)$ from Table 1.
  2. 2.
    Output the following estimator:
    $\tilde{\tau}^{(2)}_{\alpha,\beta}=\frac{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)Y_{i}+\nu_{1}}{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)+\nu_{2}}-\frac{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)Y_{i}+\nu_{3}}{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)+\nu_{4}},$
    where $\nu_{j}\sim\text{Lap}\left\{0,\frac{1}{(1-p)q_{j}\epsilon\eta}\right\}$ for $j=1,\dots,4$.
Proposition 3.
Algorithm 1 satisfies $\epsilon$-DP.
Next, we examine the asymptotic properties of $\tilde{\tau}^{(2)}_{\alpha,\beta}$. The following theorem, which addresses the consistency and rate of convergence, represents a novel contribution to the literature on causal inference under DP.
Theorem 3.2.
Under Assumptions 12 and the technical conditions in Hirano et al. (2003) included in appendix, the estimator $\tilde{\tau}^{(2)}_{\alpha,\beta}$ is consistent for $\tau_{\alpha,\beta}^{*}$. For all values of $\alpha,\beta\in[0,1]$ except $\alpha=-1,\beta=0$ or $\alpha=0,\beta=-1$, the mean squared error of $\tilde{\tau}^{(2)}_{\alpha,\beta}$ is $O\{n^{-1}+(n\epsilon Q)^{-2}\}$, where $Q=\min\{p,(1-p)q_{1},(1-p)q_{2},(1-p)q_{3},(1-p)q_{4}\}$ is a constant when $p$ and $q_{i}$ are fixed.
This theorem implies that our private estimator is optimal because its mean squared error matches the minimax risk lower bound in Lemma 1 , ensuring rate optimality among all estimators under any privacy mechanism. In comparison to past literature, Lee et al. (2019) primarily focused on the deviation of the private estimator from the non-private estimator, and their estimator remains biased for the true causal effect, even asymptotically, due to the regularization necessary for objective perturbation to produce the privatized propensity score. While Guha and Reiter (2024) demonstrated the consistency of their estimator, their asymptotic guarantees are contingent upon the choice of the number of splits $M$ in the subsample-and-aggregate algorithm, which serves as a hyperparameter of the algorithm. The convergence rate they present is also expressed in terms of the hyperparameter $M$ rather than the sample size $n$.
3.5 Variance Estimator
To estimate the variance and construct the confidence interval of the estimator, we follow the strategy of Li et al. (2018), as described below. By the law of total variance, the variance of the (non-private) estimator (7) is $\mathrm{var}(\hat{\tau}_{\alpha,\beta})=\mathrm{E}_{X}\mathrm{var}(\hat{\tau}_{\alpha,\beta})+\mathrm{var}_{X}\mathrm{E}(\hat{\tau}_{\alpha,\beta})$. They employed the argument of Imbens (2004) that individual variation $\mathrm{E}_{X}\mathrm{var}(\hat{\tau}_{\alpha,\beta})$ is typically much larger than conditional mean variation $\mathrm{var}_{X}\mathrm{E}(\hat{\tau}_{\alpha,\beta})$ to argue that the second term is negligible. Li et al. (2018, Theorem 2) further showed that, when $n$ is large, the first term $\mathrm{E}_{X}\mathrm{var}(\hat{\tau}_{\alpha,\beta})$ can be approximated by:
$V_{\alpha,\beta}=\frac{1}{nH_{\alpha,\beta}^{2}}\int h_{\alpha,\beta}(x)^{2}\left\{\frac{v_{1}(x)}{e(x)}+\frac{v_{0}(x)}{1-e(x)}\right\}f(x)\mu(dx),$ (13)
where $v_{z}(x)=\mathrm{var}\{Y(z)\mid X=x\}$ and $H_{\alpha,\beta}=\int h_{\alpha,\beta}(x)f(x)d\mu(x)$ is a normalizing constant. (13) can be estimated by the following estimator:
$\begin{split}\hat{V}_{\alpha,\beta}(D)&=\frac{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})^{2}\left\{\frac{\hat{v}_{1}(X_{i})}{\hat{e}(X_{i})}+\frac{\hat{v}_{0}(X_{i})}{1-\hat{e}(X_{i})}\right\}}{\left\{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})\right\}^{2}},\end{split}$ (14)
where $\hat{e}(X_{i})$ is the non-private estimator of the propensity score obtained by solving (2), $\hat{h}_{\alpha,\beta}(x)=\hat{e}(x)^{\alpha+1}\{1-\hat{e}(x)\}^{\beta+1}$ and $\hat{v}_{z}(x)$ is a unbiased estimator for $v_{z}(x)$. In practice, we will need a model or an additional assumption to estimate the variance $v_{z}(x)$. We will assume the homoscedastic variance $v_{0}(x)=v_{1}(x)=v$ across both groups, which enables us to estimate the variance by a simple unbiased estimator of the observed outcome. The following lemma establishes the $95\%$ confidence interval for $\tau_{\alpha,\beta}$ with $\hat{V}_{\alpha,\beta}$.
Lemma 3.
Under Assumptions 12 and the technical conditions in Hirano et al. (2003) included in appendix, we have an asymptotically valid $95\%$ confidence interval for $\tau_{\alpha,\beta}$: $\left(\hat{\tau}_{\alpha,\beta}-1.96\sqrt{\hat{V}_{\alpha,\beta}},\hat{\tau}_{\alpha,\beta}+1.96\sqrt{\hat{V}_{\alpha,\beta}}\right)$.
We approximate $\hat{V}_{\alpha,\beta}$ via the two-stage privatization as we do in Section 3.1. For the first stage, we reuse $\tilde{\theta}^{(1)}_{\alpha,\beta}$. The second stage applies the Laplace mechanism to the estimator privatized in the first stage. The following lemma states the $\ell_{1}$-sensitivity of the estimator.
Lemma 4.
Suppose we use the CBSR defined by (3) and (4) with $1\leq\alpha,\beta\leq 0$. The $\ell_{1}$-sensitivity of $\hat{V}_{\alpha,\beta}$ and $\tilde{V}^{(1)}_{\alpha,\beta}$ is $\Delta_{V,\alpha,\beta}=1/2n\eta C_{\eta,\alpha,\beta}$ where $C_{\eta,\alpha,\beta}=\min\{\eta^{\alpha+1}(1-\eta)^{\beta+1},(1-\eta)^{\alpha+1}\eta^{\beta+1}\}$.
This sensitivity is a generalization of the sensitivity of Guha and Reiter (2024). Choosing appropriate values of $\alpha$ and $\beta$ for each estimand gives us the same sensitivity. Table 2 presents the correspondence of estimands and sensitivities.
We consider the following estimator for the variance. $\tilde{V}^{(2)}_{\alpha,\beta}=\tilde{V}^{(1)}_{\alpha,\beta}+\nu_{V},$ where $\nu_{V}\sim\mathrm{Lap}(1,\Delta_{V,\alpha,\beta}/\epsilon)$ and $\Delta_{V,\alpha,\beta}$ represents the $\ell_{1}$-sensitivity of $\tilde{V}^{(1)}_{\alpha,\beta}$. The variance estimator $\tilde{V}^{(2)}_{\alpha,\beta}$ satisfies $\epsilon$-DP; however, it is possible that the Laplace mechanism produces a negative value for the variance estimator. To address this issue, we apply the following post-processing:
$\tilde{V}_{\alpha,\beta}^{*}=\begin{cases}\tilde{V}^{(2)}_{\alpha,\beta}&\text{ if }\tilde{V}^{(2)}_{\alpha,\beta}>0\\ \frac{1}{4n\eta C_{\eta,\alpha,\beta}}+\frac{1}{2\epsilon^{2}n^{2}\eta^{2}}&\text{ if }\tilde{V}_{\alpha,\beta}\leq 0.\end{cases}$ (15)
For $\tilde{V}^{(2)}_{\alpha,\beta}\leq 0$, the first term is the upper-bound of $\hat{V}_{\alpha,\beta}$, and the second term is the variance of the noise from the Laplace mechanism. When using this post-processing in a plug-in confidence interval, we obtain a conservative confidence interval with a coverage probability greater than the nominal confidence level. Algorithm 2 summarizes the process for obtaining $\tilde{V}^{*}_{\alpha,\beta}$, and Theorem 3.3 ensures the privacy guarantee of Algorithm 2.
Algorithm 2 Differentially Private Variance Estimator
0: Database $D=\{(X_{i},Z_{i},Y_{i})\}_{i=1}^{n}$, positivity bound $\eta$, privacy budget $\epsilon$, privatized propensity score $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ from Algorithm 1.
0: Differentially private variance estimator $\tilde{V}^{*}_{\alpha,\beta}$.
1: Compute the non-private estimator by plugging in the propensity score $\hat{e}(x)$ obtained by solving the optimization equation.
2: Compute the partially privatized estimator $\tilde{V}^{(1)}_{\alpha,\beta}$ by plugging in $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(x)$ for $\hat{e}(x)$ in the non-private variance estimator.
3: Apply the Laplace mechanism to privatize the variance estimator:
$\tilde{V}^{(2)}_{\alpha,\beta}=\tilde{V}^{(1)}_{\alpha,\beta}+\nu_{V}$
where $\nu_{V}\sim\text{Lap}\left(0,\frac{\Delta_{V,\alpha,\beta}}{\epsilon}\right)$, and $\Delta_{V,\alpha,\beta}$ is selected based on the estimand of interest from Table 2.
4: Apply post-processing to ensure positivity of the variance estimator:
$\tilde{V}^{*}_{\alpha,\beta}=\begin{cases}\tilde{V}^{(2)}_{\alpha,\beta}&\text{if }\tilde{V}^{(2)}_{\alpha,\beta}>0,\\ \frac{1}{4n\eta C_{\eta,\alpha,\beta}}+\frac{1}{2\epsilon^{2}n^{2}\eta^{2}}&\text{if }\tilde{V}^{(2)}_{\alpha,\beta}\leq 0.\end{cases}$
5:return  $\tilde{V}^{*}_{\alpha,\beta}$.
Theorem 3.3.
Algorithm 2 satisfies $\epsilon$-DP. Under Assumption 12 and the technical conditions in Hirano et al. (2003) included in appendix, $\tilde{V}^{*}_{\alpha,\beta}$ is consistent for $V_{\alpha,\beta}$.
Remark (Choice of hyperparameters).
Since both $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$ separately satisfy $\epsilon$-DP, to obtain $\epsilon$-DP for the joint release, we need to scale the privacy budgets accordingly, but we do not need to allocate an equal budget to both $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$. Let $0<r<1$ denote the budget proportion for $\tilde{V}^{*}_{\alpha,\beta}$. We then apply the privacy mechanisms for $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$ such that they satisfy $(1-r)\epsilon$-DP and $r\epsilon$-DP respectively. Notably, we can reuse $\tilde{\theta}^{(1)}_{\alpha,\beta}$ for privatizing $\hat{V}^{*}_{\alpha,\beta}$, therefore, it is recommended that we allocate a smaller budget on $\tilde{V}^{*}_{\alpha,\beta}$, i.e., $r<0.5$.
Specifically, our methodology involves three key hyperparameters: $p\in(0,1)$, $0<q_{j}<1$ for $j=1,\ldots,4$ with $\sum_{j=1}^{4}q_{j}=1$, and $r\in(0,1)$. The parameters $p$ and $q_{j}$ govern the privacy level of the point estimator, specifically controlling the allocation of the privacy budget between the first and second stages of privatization. When there is no prior information to guide the choice of these parameters, it is generally recommended to assign equal weights to each privatization. Specifically, we set $p=0.5$ and $q_{1}=\ldots=q_{4}=0.25$ as the default choice.
The parameter $r$ controls the fraction of the privacy budget allocated to the variance estimator. As discussed in Section 3.4, the privatization of the variance estimator benefits from the already privatized propensity score from the first stage of point estimator privatization. Consequently, it is recommended to allocate a smaller portion of the privacy budget to the variance estimator compared to the point estimator. We set $r=1/6$ to equally allocate the budget to the variance estimator and all five components of the point estimator.
Finally, the positivity bound $\eta$ affects the estimators. It enters the variance estimator and the confidence intervals directly, whereas the point estimator is less sensitive. In our illustrative analysis, we use $\eta=0.05$ as the default. For applications in practice, we recommend assessing sensitivity to alternative choices of $\eta$.
4 Simulation Studies
4.1 Setup
In this section, we evaluate the frequentist properties of our methodologies through repeated sampling under various privacy budgets. The evaluation metrics include mean squared error and relative bias of the point estimator, and the $95\%$ coverage and the interval length of the interval estimator. These metrics are formally defined as follows: mean squared error is calculated as $1/N_{\mathrm{sim}}\sum_{n=1}^{N_{\mathrm{sim}}}\left(\tau-\hat{\tau}_{n}\right)^{2}$, relative bias as $1/N_{\mathrm{sim}}\sum_{n=1}^{N_{\mathrm{sim}}}\left|\tau-\hat{\tau}_{n}\right|/\tau$, coverage as $1/N_{\mathrm{sim}}\sum_{n=1}^{N_{\mathrm{sim}}}\mathbbm{1}\left(\hat{\tau}_{n}^{l}\leq\tau\leq\hat{\tau}_{n}^{u}\right)$, and interval length as $1/N_{\mathrm{sim}}\sum_{n=1}^{N_{\mathrm{sim}}}\left(\hat{\tau}_{n}^{u}-\hat{\tau}_{n}^{l}\right)$ , where $N_{\mathrm{sim}}$ denotes the number of repeated samples, $\tau$ represents the true causal estimand, and $\hat{\tau}_{n}$, $\hat{\tau}_{n}^{l}$, and $\hat{\tau}_{n}^{u}$ denote the point estimate of the causal estimand, and the lower and upper bounds of the $95\%$ confidence interval for the causal estimand, respectively. In this study, we use $N_{\mathrm{sim}}=300$ datasets. We compare our methodology with existing approaches documented in the literature (Lee et al., 2019; Guha and Reiter, 2024). Although Lee et al. (2019) focused on $(\epsilon,\delta)$-DP, their methodology can be extended to $\epsilon$-DP easily using the Algorithm 6 of Awan and Slavković (2021).
We present simulation studies designed to provide insight into the discriminating capability of our method over existing methodologies. We generate various size of observations $N\in\{5000,10000,50000,100000\}$, each with $d=4$ covariates, denoted as $X_{i}=(X_{i1},X_{i2},X_{i3},X_{i4})$. The covariates are simulated from a multivariate normal distribution, $X_{i}\sim N\{0,(1-\rho)I+\rho J\}$, where $0\leq\rho\leq 1$, and $J$ is a $p\times p$ matrix with each entry equal to 1. We set $\rho=0.2$. We rescale $X_{i}$ such that $\|X_{i}\|_{2}\leq 1$. For each $(X_{i},Z_{i})$ pair, we simulate potential binary outcomes, $(Y_{i}(0),Y_{i}(1))$, from Bernoulli distributions with probabilities defined by $l[\mathrm{pr}\{Y(Z)=1\mid X_{i}\}]=\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+\beta_{3}X_{3}+\beta_{4}X_{4}+\gamma Z,$ where we set $(\beta_{0},\beta_{1},\beta_{2},\beta_{3},\beta_{4})=(0.15,-0.2,0.3,-0.4,0.6)$, and $\gamma$ controls the treatment effect, which we set to 1.
For the hyperparameters of our methodology, we chose $p=0.2$, $q_{1}=\ldots=q_{4}=0.25$ and $r=0.3$. For the hyperparameters of Guha and Reiter (2024)’s method, we set $M=\sqrt{N}$, following their simulation setups in (Guha and Reiter, 2024). We found that setting $M=\sqrt{N}$ produces better results than setting $M$ to a fixed value $M=100$. The supplementary material provides a discussion on the choice of hyperparameters used in our methodology.
4.2 Scenario 1: Correctly Specified Propensity Score Model
First, we consider a scenario where the true propensity is a logistic regression without non-linear transformations, and the working model is correctly specified. We adopt the same data-generating process as described in Guha and Reiter (2024). Specifically, the following propensity score model serves as the true model:
$l\{\mathrm{pr}(Z_{i}=1\mid X_{i})\}=0.1+0.8X_{i1}+2.0X_{i2}-1.0X_{i3}-1.8X_{i4},$
where $l(\cdot)$ is the logistic link function. This is a simple logistic regression model with linear predictors and no non-linear transformations. Given the well-specified nature of this model, methods proposed by Guha and Reiter (2024) and Lee et al. (2019) are expected to perform well.
Table 3 shows the simulation results for various sample sizes, $N\in\{5000,10000,50000\}$, and total privacy budgets, $\epsilon\in\{0.5,1.0,5.0\}$, which are allocated to each privatization steps based on the hyperparameter chosen. Our methodology exhibits superior performance even in this well-specified scenario. Specifically, our methodology consistently produces smaller bias and mean squared error across all scenarios. The Lee et al. (2019) method exhibits the largest bias and mean squared error, which can be attributed to the regularization bias inherent in the objective perturbation and the use of the Horvitz-Thompson estimator (Horvitz and Thompson, 1952), whereas the Guha and Reiter (2024) and our methods employ the Hájek estimator with self-normalizing weights. It is well-known that the Hájek estimator generally has a smaller variance than the Horvitz-Thompson estimator (Hirano et al., 2003). Compared to the Guha and Reiter (2024)’s method, our methodology yields smaller mean squared error and bias across all scenarios. This result highlights the critical importance of achieving good covariate balance and optimal rates.
An additional key observation is that our estimator yields smaller bias and mean squared error at a faster rate than the Guha and Reiter (2024)’s method. As shown in Table 4, the performance gap between our method and the Guha and Reiter (2024)’s method widens as $N$ increases. This improvement is due to the asymptotic properties of our estimator. While the convergence rate of our estimator decays with $N$, the rate of Guha and Reiter (2024)’s method decays with $M$, a hyperparameter of their algorithm.
4.3 Scenario 2: Mispecified Propensity Score Model
Next, we consider a scenario where the working model is misspecified as a simple generalized linear model, whereas the true propensity score model is nonlinear. This scenario is particularly relevant in practice, as propensity score models are often modeled as simple logistic regressions but are rarely correctly specified. For $i=1,\dots,N$, we generate the treatment variable $Z_{i}$ from a nonlinear propensity score model with the probability $\mathrm{pr}(Z_{i}=1\mid X_{i})$ given by
$l\{\mathrm{pr}(Z_{i}=1\mid X_{i})\}=0.1+0.4\exp(-X_{i1}/2)+X_{i2}X_{i3}-0.6\sin(X_{i1})-0.9X_{i4}^{2}.$
For inference, we adopt the following misspecified linear logistic regression model: $l\{\mathrm{pr}(Z_{i}=1\mid X_{i};\theta)\}=X_{i}^{\top}\theta$.
Table 4 presents the simulation results. Our methodology consistently produces smaller bias and mean squared error across all scenarios, as it did under the correctly specified scenario. In terms of coverage, our method yields conservative coverage probabilities always greater than $95\%$, which we expected from the construction of our private variance estimator. On the other hand, the Guha and Reiter (2024)’s method does not exhibit calibrated coverage probabilities for large $N$ and small $\epsilon$. The inferior performance of the Guha and Reiter (2024) and LGPR method for all metrics can be attributed to model misspecification. Our methodology offers greater robustness to model misspecification and improved efficiency in finite samples by ensuring good covariate balance between groups, consistent with the literature emphasizing the importance of covariate balance under model misspecification (Imai and Ratkovic, 2014; Zhao and Percival, 2017; Chan et al., 2015; Zhao, 2019).
Table 3: Evaluation metrics under the well-specified scenario for our differentially private methodology versus the method of Guha and Reiter (2024) and Lee et al. (2019) under various sample sizes $N\in\{5000,10000,50000,100000\}$ and privacy budgets $\epsilon\in\{0.5,1.0,5.0\}$. The Lee et al. (2019)’s method does not offer the interval estimator, thus the coverage and interval length are not displayed. GR and LGPM represent the method of Guha and Reiter (2024) and Lee et al. (2019), repectively.
MSE Bias Coverage Interval Length NN $\epsilon$ Our Method GR LGPM Our Method GR LGPM Our Method GR Our Method GR 0.5 $0.01425$ $0.02646$ $0.01266$ $0.41543$ $0.53978$ $0.39716$ $95.3\%$ $94.7\%$ $0.70084$ $0.65664$ 5000 1.0 $0.00378$ $0.00692$ $0.00625$ $0.21576$ $0.27788$ $0.29859$ $97.0\%$ $98.0\%$ $0.49685$ $0.46435$ 5.0 $0.00043$ $0.00067$ $0.00477$ $0.07581$ $0.09274$ $0.28739$ $97.0\%$ $100.0\%$ $0.22206$ $0.20795$ 0.5 $0.00394$ $0.01557$ $0.00556$ $0.21269$ $0.39357$ $0.29409$ $95.3\%$ $95.7\%$ $0.48965$ $0.54914$ 10000 1.0 $0.00110$ $0.00412$ $0.00447$ $0.11445$ $0.20580$ $0.27928$ $98.7\%$ $98.0\%$ $0.34867$ $0.38823$ 5.0 $0.00019$ $0.00036$ $0.00423$ $0.05023$ $0.06580$ $0.28273$ $97.7\%$ $100.0\%$ $0.15619$ $0.17357$ 0.5 $0.00016$ $0.00229$ $0.00398$ $0.04670$ $0.16421$ $0.28048$ $99.0\%$ $100.0\%$ $0.21604$ $0.36941$ 50000 1.0 $0.00007$ $0.00062$ $0.00391$ $0.02942$ $0.08585$ $0.28036$ $99.3\%$ $100.0\%$ $0.15383$ $0.26131$ 5.0 $0.00003$ $0.00007$ $0.00399$ $0.02117$ $0.03048$ $0.28413$ $96.0\%$ $100.0\%$ $0.06996$ $0.11668$ 0.5 $0.00005$ $0.00136$ $0.00383$ $0.02373$ $0.11743$ $0.27863$ $9.0\% 99$$9.7\%$ $0.15522$ $0.30914$ 100000 1.0 $0.00002$ $0.00036$ $0.00384$ $0.01679$ $0.06168$ $0.27962$ $99.7\%$ $0.0\% 0.110$520.11052 $0.21858$ 5.0 $0.00001$ $0.00004$ $0.00385$ $0.01380$ $0.02188$ $0.28008$ $99.0\%$ $100.0\%$ $0.04785$ $0.09769$
Table 4: Evaluation metrics under the misspecified scenario.
MSE Bias Coverage Interval Length NN $\epsilon$ Our Method GR LGPM Our Method GR LGPM Our Method GR Our Method GR 0.5 $0.01682$ $0.02637$ $0.06417$ $0.33168$ $0.38982$ $0.78137$ $93.7\%$ $22.0\%$ $0.69825$ $0.05990$ 5000 1.0 $0.00421$ $0.00681$ $0.05776$ $0.16577$ $0.20005$ $0.77720$ $97.0\%$ $98.0\%$ $0.49834$ $0.46435$ 5.0 $0.00030$ $0.00041$ $0.05686$ $0.04498$ $0.05331$ $0.78577$ $99.0\%$ $100.0\%$ $0.23289$ $0.20762$ 0.5 $0.00440$ $0.01555$ $0.05687$ $0.16513$ $0.28418$ $0.77599$ $94.7\%$ $19.3\%$ $0.49029$ $0.03619$ 10000 1.0 $0.00112$ $0.00391$ $0.05606$ $0.08423$ $0.14648$ $0.78001$ $97.7\%$ $98.7\%$ $0.34817$ $0.38824$ 5.0 $0.00011$ $0.00023$ $0.05629$ $0.02701$ $0.03898$ $0.78499$ $98.0\%$ $100.0\%$ $0.16101$ $0.17373$ 0.5 $0.00017$ $0.00219$ $0.05596$ $0.03481$ $0.11630$ $0.78359$ $99.3\%$ $13.7\%$ $0.21625$ $0.01217$ 50000 1.0 $0.00006$ $0.00058$ $0.05587$ $0.01983$ $0.06094$ $0.78346$ $99.7\%$ $100.0\%$ $0.15357$ $0.26129$ 5.0 $0.00002$ $0.00004$ $0.05583$ $0.01062$ $0.01701$ $0.78343$ $99.0\%$ $100.0\%$ $0.07037$ $0.11691$ 0.5 $0.00005$ $0.00134$ $0.05546$ $0.01717$ $0.08593$ $0.78102$ $99.7\%$ $12.3\%$ $0.15604$ $0.00775$ 100000 1.0 $0.00002$ $0.00035$ $0.05555$ $0.01058$ $0.04385$ $0.78172$ $99.7\%$ $0.0\% 0.110$350.11035 $0.21863$ 5.0 $0.00001$ $0.00003$ $0.05563$ $0.00738$ $0.01310$ $0.78230$ $99.7\%$ $100.0\%$ $0.05180$ $0.09784$
5 Data Analysis
In this section, we demonstrate the application of our estimators by assessing the treatment effect of a labor training program using data previously analyzed by LaLonde (1986) and Dehejia and Wahba (1999), among others. The ‘National Supported Work’ (NSW) demonstration was a randomized experiment conducted in the mid-1970s to determine if a systematic job training program could increase post-intervention income levels, measured in 1978. The data include various individual covariates, such as age, education, Black (1 if Black, 0 otherwise), Hispanic (1 if Hispanic, 0 otherwise), married (1 if married, 0 otherwise), no degree (1 if no degree, 0 otherwise), earnings in 1974, and earnings in 1975. In our study, we convert the income data into a binary outcome, representing whether an individual’s income is greater than zero or not, as a proxy for employment status.
The original NSW study included both intervention and control groups under a randomized experiment. LaLonde (1986) investigated the extent to which analyses using observational datasets as controls could replicate the unbiased results of a randomized experiment. His non-experimental estimates were derived from an observational cohort: the Panel Study of Income Dynamics (PSID). Detailed descriptions of these datasets can be found in LaLonde (1986) and Dehejia and Wahba (1999). The units in the PSID study serve as observational control data since participants in these groups did not partake in the NSW job training program. Thus, we integrated these datasets and created two comparative datasets. The first one is from the NSW study with those treated ($N=185$) and those in the control group ($N=260$), which is regarded as the experimental data. The second one combines the treated units from the NSW study with the control units from the PSID data ($N=2490$), which is regarded as observational data.
Our analyses aimed to evaluate whether our methodology could yield valid estimates under privacy considerations using the PSID dataset. Our methodologies are benchmarked against non-private baseline methods, which offer target values for our private estimates. For the non-private baseline, we employed the standard IPW and CBSR estimators.
Table 5 provides estimates of the average treatment effects and $95\%$ confidence intervals for the NSW and PSID datasets. We also examined how well our methodology balanced the covariate distributions between the treated and control groups. For the NSW data, the private CBSR estimators perform well, producing results that are quite similar to the non-private CBSR estimator. For example, the non-private CBSR method has an estimated average treatment effect of $0.112$, with a confidence interval from $0.023$ to $0.201$. The private CBSR estimators with moderate and generous privacy budgets ($\epsilon=1.0,5.0$) show similar point estimates as the non-private estimator, and their confidence intervals also indicate significant effects. The private CBSR method with a tight privacy budget ($\epsilon=0.5$) shows a slight deviation with an estimate of $0.067$. This slight variation illustrates that while privacy constraints may introduce some uncertainty, the private estimator still closely approximates the non-private results.
For the PSID data, the point estimators are reasonably accurate, but because the PSID dataset is observational, the results are not as robust as those for the NSW dataset. The non-private CBSR estimate for the PSID data is $0.164$, with a confidence interval from 0.075-0.075 to $0.403$. Under privacy constraints, the estimates, such as the private CBSR with $\epsilon=1.0$, showing $0.035$ with a confidence interval of 0.112-0.112 to $0.182$, reflect greater variability. This variability emphasizes the trade-off between privacy and precision, particularly in observational settings where the underlying data characteristics are less controlled than in experimental settings like the NSW data.
Finally, we also examine how our methodology ensures the balance of each covariate. The results and discussions are provided in the supplementary material.
Table 5: Estimated average treatment effects and confidence intervals for LaLonde (1986) data.
Results for NSW data Results for PSID data Estimator Estimate Confidence Interval Estimate Confidence Interval Non-private CBSR 0.112 $(0.023,0.201)$ 0.164 (0.075,0.403)(-0.075,0.403) Private CBSR ($\epsilon=0.5$) 0.067 (0.004,0.138)(-0.004,0.138) -0.007 (0.147,0.132)(-0.147,0.132) Private CBSR ($\epsilon=1.0$) 0.082 $(0.003,0.162)$ 0.035 (0.112,0.182)(-0.112,0.182) Private CBSR ($\epsilon=5.0$) 0.103 $(0.017,0.188)$ 0.122 (0.030,0.274)(-0.030,0.274)
6 Concluding Remarks
In this article, we proposed a differentially private causal inference methodology. Our methodology is designed for analyzing observational data while maintaining the covariate balance between treatment groups and preserving privacy guarantees. We provided privacy guarantees for the proposed estimators, as well as the asymptotic properties, and validated their performance through simulation studies and empirical analyses using real-world data. The simulation study demonstrated that our methodology outperforms existing methods under misspecifications of the propensity score model. This robustness is due to the focus on covariate balance.
A promising avenue for future research is the development of an analytical framework for unbounded variables. Our current framework is restricted to bounded variables due to the sensitivity considerations of differential privacy mechanisms. Additionally, the finite-sample performance of our estimators could potentially be improved by more carefully designing the second stage of our privatization process, allowing for a more tailored noise distribution (Awan and Slavković, 2021). Finally, because the proposed methodology targets low- and moderate-dimensional covariates, extending it to high-dimensional settings is important. In particular, the rejection-sampling framework of Awan and Rao (2024) can be computationally burdensome in high dimensions, so more efficient privatization algorithms are needed.
Appendix A Technical Proofs
A.1 Proof of Lemma 1
Proof.
As discussed in Section 2, we consider a scenario where each unit $i$ has an outcome $Y_{i}\in[0,1]$, treatment assignment $Z_{i}\in\{0,1\}$, and covariates $X_{i}\in\mathcal{X}$, where $\mathcal{X}$ is the unit ball so that $\|X_{i}\|_{2}\leq 1$, respectively. Along with the boundedness assumptions, Proposition 2 of Barber and Duchi (2014) implies the minimax risk is lower bounded $\inf_{M_{\epsilon}\in\mathcal{M}_{\epsilon}}\inf_{\hat{\tau}}\sup_{P\in\mathcal{P}}\mathrm{E}\left\{(\hat{\tau}-\tau)^{2}\right\}=\Omega\left\{n^{-1}+(\epsilon n)^{-2}\right\}$. ∎
A.2 Proof of Lemma 2
Proof.
By Equations (3), (4) and the chain rule, we can compute the gradient of $S_{n,\alpha,\beta}(\theta;D)$ as
$\nabla_{\theta}S_{n,\alpha,\beta}(\theta,D)$ $=\frac{\partial e}{\partial\theta}\frac{\partial}{\partial e}S_{n,\alpha,\beta}(\theta,D)$
$=\sum_{i=1}^{n}\frac{\partial e}{\partial\theta}(Z_{i}-e_{\theta}(X_{i}))G^{\prime\prime}(e_{\theta}(X_{i}))$
$=\sum_{i=1}^{n}\frac{\partial e}{\partial\theta}(Z_{i}-e_{\theta}(X_{i}))e_{\theta}(X_{i})^{\alpha-1}(1-e_{\theta}(X_{i}))^{\beta-1}$
$=\sum_{i=1}^{n}\left\{(Z_{i}-e_{\theta}(X_{i}))e_{\theta}(X_{i})^{\alpha}(1-e_{\theta}(X_{i}))^{\beta}\right\}(\phi_{1}(X_{i}),\ldots,\phi_{d}(X_{i}))^{\top}.$
Also, the $\ell_{2}$-sensitivity is:
$\Delta_{\theta,\alpha,\beta}$ $=\|\nabla S_{n,\alpha,\beta}(\theta,D)-\nabla S_{n,\alpha,\beta}(\theta,D^{\prime})\|_{2}$
$\leq 2\|\left\{(Z_{i}-e_{\theta}(X_{i}))e_{\theta}(X_{i})^{\alpha}(1-e_{\theta}(X_{i}))^{\beta}\right\}(\phi_{1}(X_{i}),\ldots,\phi_{d}(X_{i}))^{\top}\|_{2}$
$\leq 2C_{\phi}|(Z_{i}-e_{\theta}(X_{i}))e_{\theta}(X_{i})^{\alpha}(1-e_{\theta}(X_{i}))^{\beta}|$
A.3 Proof of Theorem 3.1
By Zhao (2019, Theorem 1), for $j=1,\ldots,d$, we have
$\delta_{n}(\hat{\theta}_{\alpha,\beta})=\sum_{i=1}^{n}\left\{Z_{i}-(1-Z_{i})\right\}w_{\hat{\theta}_{\alpha,\beta}}(X_{i},Z_{i})\phi_{j}(X_{i})=0,$
for the non-private estimator $\hat{\theta}_{\alpha,\beta}$. From the proof of Theorem 3.2, we have $\tilde{\theta}^{(1)}_{\alpha,\beta}-\hat{\theta}_{\alpha,\beta}=O_{p}(1/n)$. Noting that $w_{\hat{\theta}_{\alpha,\beta}}(X_{i},Z_{i})$ is a function of propensity score, the continuous mapping theorem implies $\delta_{n}(\tilde{\theta}^{(1)}_{\alpha,\beta})$ converges zero in probability. To show that $\delta_{n}(\tilde{\theta}^{(1)}_{\alpha,\beta})=O_{p}(1/n)$, it suffices to show that the derivative of $\delta_{n}(\theta)$ is bounded by some constant. As discussed in the proof of Theorem 3.2, the derivative of $w(\theta)$ is bounded under Assumption 1, and the same holds for the derivative of $\delta_{n}(\theta)$. By applying the mean value theorem, we have $\delta_{n}(\tilde{\theta}^{(1)}_{\alpha,\beta})=O_{p}(1/n)$.
A.4 Proof of Proposition 3
Proof.
Releasing $\tilde{\theta}^{(1)}_{\alpha,\beta}$ achieves $p\epsilon$-DP due to the guarantee of the KNG mechanism with $\ell_{2}$ norm (Proposition 2). Noting that the $\ell_{1}$-sensitivity of each component of (12) is $1/\eta$, Proposition 1 and the composition property implies that releasing four private quantities (two numerators and two denominators) satisfies $\sum_{j=1}^{4}(1-p)q_{j}\epsilon=(1-p)\epsilon$-DP. Finally, by composition, releasing $\tilde{\tau}^{(2)}_{\alpha,\beta}$ satisfies $p\epsilon+(1-p)\epsilon=\epsilon$-DP. ∎
A.5 Proof of Theorem 3.2
We first introduce the technical assumptions in Hirano et al. (2003).
Assumption 3.
The support of $X$ is a Cartesian product of compact intervals. The density of $X$ is bounded, and bounded away from 0.
Assumption 4.
The second moments of $Y(0)$ and $Y(1)$ exist and $g(X,0)=\mathrm{E}\{Y(0)\mid X\}$ and $g(X,1)=\mathrm{E}\{Y(1)\mid X\}$ are continuously differentiable.
Assumption 5.
The propensity score $e(X)=p(Z=1|X)$ is continuously differentiable of order $s\geq 7d$ where $d$ is the dimension of $X$.
Assumption 6.
The propensity score is specified by the nonparametric sieve logistic regression that uses a power series with $m=n^{a}$ for some $\frac{1}{4(s/d-1)}<a<\frac{1}{9}$.
Proof.
First, the WATE and its estimators can be decomposed into the sum of two parts, i.e., the term relevant to $Y(1)$ and the other term relevant to $Y(0)$. For simplicity, we only consider the first term of the estimands and estimators throughout the proof, and the same proof techniques can be applied to the second term. We slightly abuse the notations to represent the first term of the estimands and estimators by $\tilde{\tau}^{(2)}_{\alpha,\beta}$, $\tilde{\tau}^{(1)}_{\alpha,\beta}$, $\hat{\tau}_{\alpha,\beta}$ and $\tau_{\alpha,\beta}$. We also suppress the subscript α,β for simplicity. Consider the following quantities: $\xi_{1}=\|\hat{\tau}-\tau\|$, $\xi_{2}=\|\tilde{\tau}^{(1)}-\tau\|$, $\xi_{3}=\|\tilde{\tau}^{(2)}-\tau\|,$ where $\tau$, $\hat{\tau}$, $\tilde{\tau}^{(1)}$, and $\tilde{\tau}^{(2)}$ represent the estimand, the non-private estimator, the private estimator after the first stage, and the final private estimator, respectively. We will study the asymptotic behavior of each component.
The convergence of the first component $\xi_{1}$ can be proved by applying the proof of Hirano et al. (2003); Zhao (2019). Under Assumptions 16, Hirano et al. (2003); Zhao (2019) showed that $\hat{\tau}$ is a semiparametric efficient estimator for $\tau$ with the convergence rate $O_{p}(1/\sqrt{n})$. Thus $\xi_{1}$ converges to zero in probability and $\xi_{1}=O_{p}(1/\sqrt{n})$.
[…]
where $\tau^{\prime}(w_{0})$ is the derivative of $\tau(w)$ evaluated at $w_{0}$. Similarly, by the mean value theorem, for some $\theta_{0}$, we have
[…]
Under Assumption 1, $\tau^{\prime}(w_{0})$ and $w^{\prime}(\theta_{0})$ are bounded. Additionally, according to Reimherr and Awan (2019, Theorem 3.2), the KNG mechanism ensures that, if the loss function is a twice differentiable convex loss function and the Hessian has strictly positive eigenvalues, we have $\tilde{\theta}-\hat{\theta}=O_{p}(1/np\epsilon).$ Zhao (2019, Proposition 2) showed this condition holds for $\alpha=-1,\beta=0$ or $\alpha=0,\beta=-1$ in the CBSR framework. Therefore, putting these together, we have
$\xi_{2}=\tau(\tilde{w})-\tau(\hat{w})+\xi_{1}=O_{p}(1/np\epsilon+1/\sqrt{n}).$
Finally, consider $\xi_{3}$. Given that $\tilde{\tau}^{(1)}=\tau(\tilde{w})=\tau+O_{p}(1/np\epsilon+1/\sqrt{n})$, we have
$\tilde{\tau}^{(2)}$ $=\frac{\sum_{i=1}^{n}\tilde{w}_{i}Z_{i}Y_{i}+\nu_{1}}{\sum_{i=1}^{n}\tilde{w}_{i}Z_{i}+\nu_{2}}$
$=\frac{\tau\sum_{i=1}^{n}\tilde{w}_{i}Z_{i}+O_{p}(1/np\epsilon+1/\sqrt{n})\sum_{i=1}^{n}\tilde{w}_{i}Z_{i}+\nu_{1}}{\sum_{i=1}^{n}\tilde{w}_{i}Z_{i}+\nu_{2}}$
$=\frac{\tau a_{n}+O_{p}(1/np\epsilon+1/\sqrt{n})a_{n}+\nu_{1}}{a_{n}+\nu_{2}}$
$=\frac{\tau+O_{p}(1/np\epsilon+1/\sqrt{n})+\nu_{1}/a_{n}}{1+\nu_{2}/a_{n}}$
$=\{\tau+O_{p}(1/np\epsilon+1/\sqrt{n})+\nu_{1}/a_{n}\}\{1-\nu_{2}/a_{n}+(\nu_{2}/a_{n})^{2}-\cdots\}$
$=\tau+O_{p}(1/np\epsilon+1/\sqrt{n})+\nu_{1}/a_{n}-\tau\nu_{2}/a_{n}-\nu_{2}/a_{n}O_{p}(1/np\epsilon+1/\sqrt{n})-\nu_{1}\nu_{2}/a_{n}^{2},$
where the third line follows from $a_{n}=\sum_{i=1}^{n}\tilde{w}_{i}Z_{i}$, the fifth line follows from the Taylor expansion, representing the higher-order terms that are asymptotically negligible. The last line also follows because the higher-order terms are asymptotically negligible. Under Assumption 1, we have $a_{n}=O_{p}(n)$. Putting all of these together, we have
$\xi_{3}=\tilde{\tau}^{(2)}-\tau$ $=O_{p}\left\{1/np\epsilon+1/n(1-p)q_{1}\epsilon+1/n(1-p)q_{2}\epsilon+1/\sqrt{n}\right\}$
$=O_{p}\left[1/\sqrt{n}+\left\{\min(np\epsilon,n(1-p)q_{1}\epsilon,n(1-p)q_{2}\epsilon\right\}^{-1}\right]$
Applying the same procedure to the second term of WATE, we obtain the desired result. ∎
A.6 Proof of Lemma 4
Proof.
Here, we prove the sensitivity of $\hat{V}_{\alpha,\beta}$. The sensitivity of $\tilde{V}^{(1)}_{\alpha,\beta}$ can be proved in the same way. For any $D\in\mathcal{D}_{n}$, we have
$\hat{V}_{\alpha,\beta}(D)$ $=\frac{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})^{2}\left\{\frac{\hat{v}_{1}(X_{i})}{\hat{e}(X_{i})}+\frac{\hat{v}_{0}(X_{i})}{1-\hat{e}(X_{i})}\right\}}{\left\{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})\right\}^{2}}$
$\leq\frac{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})\left\{\frac{\hat{v}_{1}(X_{i})}{\hat{e}(X_{i})}+\frac{\hat{v}_{0}(X_{i})}{1-\hat{e}(X_{i})}\right\}}{\left\{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})\right\}^{2}}$
$\leq\frac{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})\left(\frac{1}{4}+\frac{1-\eta}{4\eta}\right)}{\left\{\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})\right\}^{2}}$
$\leq\frac{1}{4\eta\sum_{i=1}^{n}\hat{h}_{\alpha,\beta}(X_{i})},$
where the first line follows from $0\leq\hat{h}_{\alpha,\beta}(X_{i})\leq 1$ and the third line follows from Assumption 1 and $v_{z}(X_{1})\leq\frac{1}{4}$ for $z=0,1$. The variance term $v_{z}(X_{1})$ takes the maximum value $\frac{1}{4}$ when the support of $Y(z)$ is $\{0,1\}$. By taking the derivative of $h_{\alpha,\beta}(e)=e^{\alpha+1}(1-e)^{\beta+1}$ with respect to $e$, it takes its minimum value when $e=\frac{\alpha+1}{\alpha+\beta+1}$, or $\eta$ if $e=\frac{\alpha+1}{\alpha+\beta+1}<\eta$. For both cases, the minimum is $\min\{\eta^{\alpha+1}(1-\eta)^{\beta+1},(1-\eta)^{\alpha+1}\eta^{\beta+1}\}$. Then, we have
$\hat{V}_{\alpha,\beta}(D)\leq\frac{1}{4n\eta C_{\eta,\alpha,\beta}},$
where $C_{\eta,\alpha,\beta}=\min\{\eta^{\alpha+1}(1-\eta)^{\beta+1},(1-\eta)^{\alpha+1}\eta^{\beta+1}\}$. Thus, the $\ell_{1}$-sensitivity is:
$\Delta_{V,\alpha,\beta}=|\hat{V}_{\alpha,\beta}(D)-\hat{V}_{\alpha,\beta}(D^{\prime})|\leq|\hat{V}_{\alpha,\beta}(D)|+|\hat{V}_{\alpha,\beta}(D^{\prime})|=\frac{1}{2n\eta C_{\eta,\alpha,\beta}}.$
A.7 Proof of Lemma 3
Proof.
Let
$\hat{\mu}_{1}=\frac{\sum_{i=1}^{n}Z_{i}w_{\hat{\theta}_{\alpha,\beta}}Y_{i}}{\sum_{i=1}^{n}Z_{i}w_{\hat{\theta}_{\alpha,\beta}}}=\frac{\bar{\phi}_{1}}{\bar{\omega}_{1}},$
where $\bar{\phi}_{1}=\sum_{i=1}^{n}\phi_{1i}/n$ and $\bar{\omega}_{1}=\sum_{i=1}^{n}\omega_{1i}/n$ with $\phi_{1i}=Z_{i}w_{\hat{\theta}_{\alpha,\beta}}Y_{i}$ and $\omega_{1i}=Z_{i}w_{\hat{\theta}_{\alpha,\beta}}$. Further, let $\mu_{1}=\mathrm{E}(\phi_{1i})$ and $\nu_{1}=\mathrm{E}(\omega_{1i})$. By the first-order Taylor expansion of $\hat{\mu}_{1}$ around $\mu_{1}$ and $\nu_{1}$, we have
$\hat{\mu}_{1}=\frac{\mu_{1}}{\nu_{1}}-\frac{1}{\nu_{1}}(\bar{\phi_{1}}-\nu_{1})-\frac{1}{\nu_{1}^{2}}(\bar{\omega}_{1}-\nu_{1})+o_{p}(1/\sqrt{n}).$
We repeat the same process for $\hat{\mu}_{0}=\frac{\sum_{i=1}^{n}(1-Z_{i})w_{\hat{\theta}_{\alpha,\beta}}Y_{i}}{\sum_{i=1}^{n}(1-Z_{i})w_{\hat{\theta}_{\alpha,\beta}}}$. Then, we have
$\hat{\tau}_{\alpha,\beta}-\tau$
$=\left(\frac{\mu_{1}}{\nu_{1}}-\mu_{1}\right)-\left(\frac{\mu_{0}}{\nu_{0}}-\mu_{0}\right)+\frac{1}{\nu_{1}}(\bar{\phi}_{1}-\mu_{1})-\frac{1}{\nu_{1}^{2}}(\bar{\omega}_{1}-\nu_{1})-\frac{1}{\nu_{0}}(\bar{\phi}_{0}-\mu_{0})+\frac{1}{\nu_{0}^{2}}(\bar{\omega}_{0}-\nu_{0})+o_{p}(1/\sqrt{n}).$
Under Assumptions 16, Hirano et al. (2003); Zhao (2019) demonstrates that
$\left(\frac{\mu_{1}}{\nu_{1}}-\mu_{1}\right)-\left(\frac{\mu_{0}}{\nu_{0}}-\mu_{0}\right)=o_{p}(1/\sqrt{n}).$
Therefore, we have
$\sqrt{n}(\hat{\tau}_{\alpha,\beta}-\tau)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{\phi_{1i}-\mu_{1}}{\nu_{1}}-\frac{\omega_{1i}-\nu_{1}}{\nu_{1}^{2}}-\frac{\phi_{0i}-\mu_{0}}{\nu_{0}}+\frac{\omega_{0i}-\nu_{0}}{\nu_{0}^{2}}\right)+o_{p}(1).$
By the central limit theorem, we have
$\sqrt{n}(\hat{\tau}_{\alpha,\beta}-\tau)\to\mathrm{N}(0,\sigma^{2}),$
where $\sigma^{2}=\mathrm{var}(\psi_{i})$ with $\psi_{i}=\frac{\phi_{1i}-\mu_{1}}{\nu_{1}}-\frac{\omega_{1i}-\nu_{1}}{\nu_{1}^{2}}-\frac{\phi_{0i}-\mu_{0}}{\nu_{0}}+\frac{\omega_{0i}-\nu_{0}}{\nu_{0}^{2}}$. Asymptotically, $\mathrm{var}(\hat{\tau}_{\alpha,\beta})=\mathrm{var}(\psi_{i})/n$ and $\mathrm{var}(\hat{\tau}_{\alpha,\beta})$ is approximated by $\hat{V}_{\alpha,\beta}$. Thus, we obtain the desired result. ∎
A.8 Proof of Theorem 3.3
Proof.
By applying the Laplace mechanism with the sensitivity given by Lemma 4, $\tilde{V}_{\alpha,\beta}$ satisfies $\epsilon$-DP. By the post-processing property, $\tilde{V}_{\alpha,\beta}^{*}$ also satisfies $\epsilon$-DP.
Next, we consider the consistency of $\tilde{V}_{\alpha,\beta}^{*}$ for $V_{\alpha,\beta}$. First, we consider the consistency of $\tilde{V}^{(2)}_{\alpha,\beta}$. Notice that
$\|V_{\alpha,\beta}-\tilde{V}^{(2)}_{\alpha,\beta}\|$
$=$ $\|V_{\alpha,\beta}-\hat{V}_{\alpha,\beta}+\hat{V}_{\alpha,\beta}-\tilde{V}^{(1)}_{\alpha,\beta}+\tilde{V}^{(1)}_{\alpha,\beta}-\tilde{V}^{(2)}_{\alpha,\beta}\|$
$\leq$ $\underbrace{\|V_{\alpha,\beta}-\hat{V}_{\alpha,\beta}\|}_{\xi_{1}}+\underbrace{\|\hat{V}_{\alpha,\beta}-\tilde{V}^{(1)}_{\alpha,\beta}\|}_{\xi_{2}}+\underbrace{\|\tilde{V}^{(1)}_{\alpha,\beta}-\tilde{V}^{(2)}_{\alpha,\beta}\|}_{\xi_{3}}.$
By (Li et al., 2018, Theorem 2), the first term $\xi_{1}$ converges to zero. For $\xi_{2}$, we have $\tilde{\theta}-\hat{\theta}=O_{p}(1/np\epsilon)$ from the proof of Theorem 3.2. As the propensity score $e(x)$ are continuous and has bounded derivative, $\hat{e}(x)-e_{\tilde{\theta}^{(1)}}=O_{p}(1/np\epsilon)$ as well by the first order taylor expansion. For the third term $\xi_{3}$, $\tilde{V}_{\alpha,\beta}=\hat{V}_{\alpha,\beta}+\nu_{V},$ where $\nu_{V}\sim\mathrm{Lap}(1,\Delta_{V,\alpha,\beta}/\epsilon)$ and $\Delta_{V,\alpha,\beta}=\frac{1}{2n\eta C_{\eta,\alpha,\beta}}$ by Lemma 4. The Laplace noise is asymptotically negligible and decays much faster than the rate of $n^{-1/2}$, thus $\xi_{3}$ converges to zero as well, which proves the consistency of $\tilde{V}_{\alpha,\beta}^{(2)}$. Finally, the clamping post-processing (15) only serves to obtain the valid variance estimator by projecting the negative estimator to the positive upper bound of the estimator. Since the convergence rate of $\hat{V}_{\alpha,\beta}$, $O_{p}(n^{-1/2})$, is slower than the rate of the Laplace noise $\nu_{V}$, $O_{p}(n^{-1})$, the post-processing will not be applied when $n$ is large enough as the Laplace noise $\nu_{V}$ tends to zero. ∎
Appendix B Sieve Logistic Regression
We adopt the logistic regression with a polynomial series of the covariates for the propensity score estimation. This model can be interpreted as the sieve estimator (Geman and Hwang, 1982). The sieve estimators are a class of non-parametric estimators that use progressively more complex models to estimate an unknown function as more data becomes available.
The logistic regression is a widely used method for modeling the propensity score, assuming a linear dependence of the covariates $X=(X_{1},X_{2},\dots,X_{p})$ on the log-odds. While this parametric approach is computationally efficient and interpretable, it can be too restrictive in cases where the relationship between $X$ and $Y$ is nonlinear or complex. To address these limitations, sieve logistic regression provides a flexible extension by allowing for a more general specification of the relationship between $X$ and the log-odds function. In sieve logistic regression, the log-odds function is approximated using a sequence of basis functions $\{\phi_{j}(X)\}_{j=1}^{K}$ that increase in complexity as the sample size $n$ grows. This gives the model the capacity to capture nonlinear relationships without imposing a strong parametric form. Specifically, the sieve logistic regression model can be written as $\eta(X)=\sum_{j=1}^{K}\beta_{j}\phi_{j}(X),$ where $\phi_{j}(X)$ are basis functions (e.g., polynomials, splines, or Fourier expansions) chosen to flexibly approximate the true underlying relationship. The number of basis functions $K$ grows with the sample size $n$, making the model more flexible while preserving consistency.
An important choice in sieve logistic regression is the type of basis functions used. One common approach is to use orthogonalized basis functions to ensure numerical stability and efficient computation. Specifically, let $\phi(X)=(\phi_{1}(X),\phi_{2}(X),\dots,\phi_{K}(X))^{\top}$. The sieve methods approximate an arbitrary function $f:\mathbb{R}^{r}\rightarrow\mathbb{R}$ by $\theta^{\top}\phi(x)$. Because $\theta^{\top}A_{K}^{-1}A_{K}\phi(x)$, we can also use $R_{K}(x)=A_{K}\phi(x)$ as the basis of approximation. By choosing $A_{K}$ appropriately, we obtain a system of orthogonal (with respect to some weight function) functions. Specifically, we choose $A_{K}$ so that $E\{R_{K}(X)R_{K}(X)^{\top}\}=I_{K}$.
The sieve approach is particularly appealing when modeling the propensity score. Hirano et al. (2003) studied the semiparametric efficiency of the IPW estimator when the dimension of the regressors $\phi(x)$ is allowed to increase as the sample size $n$ grows. Their renowned results claim that this sieve IPW estimator is semiparametrically efficient for estimating the WATE. Zhao (2019) showed that the semiparametric efficiency still holds if the Bernoulli likelihood, the loss function that Hirano et al. (2003) used to estimate the propensity score, is replaced by the Beta family of scoring rules $G_{\alpha,\beta}$, $1\leq\alpha,\beta\leq 0$ or essentially any strongly concave scoring rule.
Specifically, using sieve logistic regression, where the series expansion of the log-odds function captures the complexities of treatment assignment mechanisms, as demonstrated by Hirano et al. (2003). By employing the sieve logistic series estimator, the propensity score is approximated as:
$e(X)=\frac{e^{R_{K}(X)^{\prime}\theta_{K}}}{1+e^{R_{K}(X)^{\prime}\theta_{K}}},$
where $R_{K}(X)$ is a vector of orthogonalized basis functions and $\theta_{K}$ is the corresponding set of estimated coefficients. As the number of basis functions $K$ increases with the sample size, the model becomes more flexible, allowing it to adapt to the complexity of the data while maintaining asymptotic properties, including consistency and efficiency.
Appendix C Privacy-Aware Rejection Sampler for the KNG Mechanism
In this section, we state the sampling algorithm we use in Algorithm 1 for private propensity score and a required lemma from Awan and Rao (2024). When the score function $S_{n,\alpha,\beta}$ for the causal estimand of interest is strongly concave, we adopt the privacy-aware rejection sampling technique of Awan and Rao (2024) to obtain a private parameter $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from (10). This algorithm is designed to obtain the exact samples from the unnormalized density $\exp(S_{D}(x))$, not relying on approximate sampling techniques such as Markov Chain Monte Carlo. For this method, we assume that for the unnormalized target density $\pi_{D}$, we have a normalized density $U_{D}(x)$ as well as a constant and $c_{U,D}$ such that for all $x$:
$\tilde{\pi}_{D}(x)\leq c_{U,D}U_{D}(x).$
Sampling from (10) falls into this type of sampling problem. Under these conditions, the privacy-aware rejection sampling algorithm is provided as in Algorithm 3.
Algorithm 3 Privacy-aware rejection sampling via squeeze functions (Awan and Rao, 2024, Algorithm 1)
Input: $\tilde{\pi}$, $U$, and $c_{U}$, such that $\tilde{\pi}(x)\leq c_{U}U(x)$ for all $x$. Output: $X_{s}$
  1. 1.
    Set anyAccepted = FALSE
  2. 2.
    Sample $X\sim U(x)$
  3. 3.
    Sample $Y\sim\text{Unif}(0,1)$
  4. 4.
    if $Y\leq\frac{\tilde{\pi}(X)}{c_{U}U(X)}$ and anyAccepted = FALSE then
  5. 5.
    Return $X_{s}$
  6. 6.
    end if
The following lemma from Awan and Rao (2024, Lemma 33) establishes the upper bound for KNG.
Lemma 5.
Let $\tilde{\pi}(x)=\exp(-\|\nabla S_{D}(x)\|^{2})$ be the unnormalized target density, where $S_{D}:\mathbb{R}^{d}\to\mathbb{R}$ is twice-differentiable and $\lambda$-strongly convex. Call $x^{*}_{D}:=\arg\min_{x}S_{D}(x)$. Write $\psi_{d}(x;m,s)$ to denote the pdf of a $d$-dimensional K-norm distribution with location $m$, scale $s$, and $\ell_{2}$ norm. Denote $\text{Vol}_{d}(\ell_{2})=\frac{2^{d}\Gamma(1+1/2)}{\Gamma(1+d/2)}$ as the volume of the unit $\ell_{2}$ ball in $\mathbb{R}^{d}$. Then, for all $x$,
$\exp(-\|\nabla S_{D}(x)\|^{2})\leq(d!)\lambda^{-d}\text{Vol}_{d}(\ell_{2})\psi_{d}(x;x^{*}_{D},1/\lambda).$
The convexity parameter $\lambda$ can be obtained from Zhao (2019). Specifically, given $\mathrm{E}(\nabla_{\theta}g_{\theta}(X))=E\{\phi(X)\phi(X)^{\top}\}=I_{d}$ from the construction of the sieve regression for $g_{\theta}(X)=\theta^{\top}\phi(X)$, it suffices to consider the following derivatives for $S_{n,\alpha,\beta}$: $\frac{d}{dg}S_{n,\alpha,\beta}(l^{-1}(g),1)=(1-e)G^{\prime\prime}(e)(l^{-1})^{\prime}(g)=e^{\alpha}(1-e)^{\beta+1}$, $\frac{d}{dg}S_{n,\alpha,\beta}(l^{-1}(g),0)=-eG^{\prime\prime}(e)(l^{-1})^{\prime}(g)=-e^{\alpha+1}(1-e)^{\beta}$, $\frac{d^{2}}{dg^{2}}S_{n,\alpha,\beta}(l^{-1}(g),1)=\alpha e^{\alpha}(1-e)^{\beta+2}-(\beta+1)e^{\alpha+1}(1-e)^{\beta+1}$, and $\frac{d^{2}}{dg^{2}}S_{n,\alpha,\beta}(l^{-1}(g),0)=-(\alpha+1)e^{\alpha+1}(1-e)^{\beta+1}+\beta e^{\alpha+2}(1-e)^{\beta}.$ Therefore, for the estimation of ATE with $\alpha=\beta=-1$ for example, we have $\lambda=\frac{n\eta}{1-\eta}$ by Assumption 1.
Remark.
Awan and Rao (2024) also established a lower bound for KNG which can be used to ensure that the runtime does not depend on the database, at the cost of slower computation time. In our setting, we assume that the runtime is not available to the statistician.
Appendix D Simulation details
According to Guha and Reiter (2024), we should choose the smallest feasible $M$ for their method, ensuring that the standard deviation of the Laplace noise from the subsampling and aggregation process remains significantly smaller than the sensitivity of the estimated variance of the WATE estimate obtained from the full data (Guha and Reiter, 2024), suggesting a fixed small value of $M$, e.g., $M=100$. However, we found that their method is sensitive to the choice of $M$ and works even better when $M$ grows with $n$. This observation aligns with their asymptotic results, where their estimator converges with $O_{p}(M^{-1})$. Therefore, we chose $M=\sqrt{N}$ instead of a fixed $M$, following their choice in their simulation.
Appendix E Addtional analyses
E.1 Covariate Balance Check for LaLonde (1986) Data under Privacy
Table 6 presents the first moment balancing measure, $\delta_{n}=\sum_{i=1}^{n}\{Z_{i}-(1-Z_{i})\}w(X_{i},Z_{i})X_{i}$, illustrating how effectively our methodology balances the covariates. $\delta_{n}=0$ indicates the perfect balance. The “Unadjusted Balance” row shows the balancing measure calculated by setting $w_{i}=1/n_{z}$, where $n_{z}=\sum_{i=1}^{n}\mathbbm{1}(Z_{i}=z)$, assigning equal weight to each individual receiving the same treatment. The metrics demonstrate that our methodology significantly balances covariates between the treatment and control groups, even under privacy constraints, as indicated by the small values of $\delta_{n}$ across different covariates. For example, age and education show very small absolute $\delta_{n}$ values, such as $0.006$ and $0.007$ for the non-private CBSR, suggesting minimal differences in these covariates between groups. The covariate balance metrics remain reasonably stable under privacy constraints, though some minor deviations are observed for the Married covariate, where the original data is not well-balanced. This highlights that, while privacy protections may introduce slight imbalances, the overall methodology still achieves strong covariate balance across most variables.
Table 6: Covariate balance metrics for PSID data. The unadjusted balance is calculated by letting $w_{i}=1/n_{z}$ where $n_{z}=\sum_{i=1}^{n}\mathbbm{1}(Z_{i}=z)$, assigning equal weight to each individual receiving the same treatment.
Balance metric: $\delta_{n}=\sum_{i=1}^{n}{Z_{i}-(1-Z_{i})}w(X_{i},Z_{i})X_{i}$ Estimator Age Education Black Hispanic Married No degree Income (1974) Income (1975) Unadjusted Balance -0.5954 -0.6274 -0.1829 -0.0281 -0.8472 -0.2476 -0.1405 -0.1210 Non-private CBSR 0.0060 0.0070 0.0062 0.0002 0.0067 0.0052 0.0010 0.0008 Private CBSR ($\epsilon=0.5$) -0.0135 -0.0103 -0.0261 0.0038 -0.2679 0.0682 -0.0146 -0.0096 Private CBSR ($\epsilon=1.0$) -0.0135 -0.0114 -0.0215 0.0038 -0.2671 0.0750 -0.0148 -0.0094 Private CBSR ($\epsilon=5.0$) -0.0130 -0.0110 -0.0196 0.0033 -0.2648 0.0720 -0.0146 -0.0094